If 4 times the reciprocal of a number is 3 more than 5/2 times the reciprocal of that number, find the number.
Answer by mananth(16946)

If 4 times the reciprocal of a number is 3 more than 5/2 times the reciprocal of that number, find the number.
let the number be x
4 times reciprocal = 4/x
(5/2x) +3
4/x = (5/2x)+3
4/x-5/2x=3
(8-5)/2x=3
3=6x
x=1/2